# Adaptive FEM with explicit time integration for the wave equation

### v1

@@ 3.3 Local time-stepping
Here we briefly recall the leapfrog (LF) based local time-stepping (LF-LTS) method used for time integration during the adaptive algorithm to circumvent the bottleneck otherwise caused by local refinement on standard explicit time-stepping methods. The original LF-LTS method for the numerical solution of the second-order wave equations was proposed for homogeneous right-hand sides in [DiazGrote:09:article:Energy-Conserving] and for inhomogeneous right-hand sides in [GroteMitkova:10:article:Explicit-Local, Sect. 4.1]. Inside the locally refined region, the latter applies standard LF time-marching with a smaller time-step, which also implies evaluating the inhomogeneous right-hand side at all intermediate times. Although we focus on second-order LTS-LF here, we note that higher-order versions are available [DiazGrote:09:article:Energy-Conserving]. For $f=0$ they also conserve (a discrete version of) the energy. Optimal convergence rates for the LF-LTS method from [DiazGrote:09:article:Energy-Conserving] with $p$ local time steps were derived for a conforming FEM discretization, albeit under a sub-optimal CFL condition where $\Delta t$ in fact depends on the smallest elements in the mesh [GroteMehlinSauter:18:article:Convergence].
To prove optimal $L^{2}$ convergence rates under a CFL condition independent of $p$, a stabilized algorithm LF-LTS$(\nu)$ was recently introduced in [GroteMichelSauter:21:article:Stabilized], and also independently in [CarleHochbruck:22:article:Error-Analysis]. Here, $\nu\geq 0$ denotes a small stabilization parameter; typically, we set $\nu=0.01$. Stability and convergence for the stabilized version LF-LTS$(\nu)$ were proved for homogeneous right-hand sides in [GroteMichelSauter:21:article:Stabilized] under a CFL condition independent of the coarse/fine mesh ratio.
Stabilization was also recently introduced into the original LTS-LF$(0)$ algorithm from [GroteMitkova:10:article:Explicit-Local] for nonzero source terms, which again led to optimal $L^{2}$-convergence rates under a CFL condition independent of the coarse-to-fine mesh ratio [GroteMichelSauter:24:ExplicitLocalTimeSteppingInhomogeneous]. Similar optimal $L^{2}$-convergence rates for a somewhat simpler ”split-LFC” variant, which omits intermediate source evaluations inside the refined region, were recently proved in [CarleHochbruck:24:article:Error].
@@
The stabilized version is denoted by LF-LTS$(\nu)$; for $\nu=0$, the original LF-LTS$(0)$ method is recovered.
@@ 3.4 Space-time adaptive algorithm
Starting from a fixed ”coarse” FE mesh, $\mathcal{M}_{0}$, which discretizes the computational domain $\Omega$, the mesh is first adapted to the initial conditions $u_{0},v_{0}$ see algorithm 2. This yields the initial mesh $\mathcal{M}_{1}$ used for time integration. Now, during each time step, we proceed as follows. First, we tentatively set $\mathcal{M}_{n+1}$ to the current mesh $\mathcal{M}_{n}$ and compute $U^{n+1}$ using the LTS-LF method inside locally refined regions and the standard LF method elsewhere. Next, we evaluate the elliptic error indicators $\varepsilon_{0}$ and $\varepsilon_{1}$ in (3.6) from $U^{n+1}$ and refine $\mathcal{M}_{n+1}$ accordingly, if needed. We then transfer $U^{n-1}$ and $U^{n}$ to the new finite element space $\mathbbmss{V}_{n+1}$, associated with $\mathcal{M}_{n+1}$ via interpolation or $\operatorname{L}^{2}$-projection, and repeat the computation of $U^{n+1}$. This process – refining $\mathcal{M}_{n+1}$ and recomputing $U^{n+1}$ – is iterated until $\varepsilon_{0}$ or $\varepsilon_{1}$ falls below the threshold $\operatorname{tol}_{H}/{N}$. At that point, $U^{n+1}$ is accepted on the current mesh $\mathcal{M}_{n+1}$. Now, the algorithm enters a coarsening step: for each $K\in\mathcal{M}_{n+1}$, we evaluate the pre-indicators $\beta_{K}$ from (3.18) and mark those elements with $\beta_{K}\leq\operatorname{tol}_{C}$. After coarsening, both $U^{n}$ and $U^{n+1}$ are transfered to the new, coarsened space $\mathbbmss{V}_{n+1}$. The entire procedure is repeated until the final time $T$ is reached.
@@
1:initial conditions $u_{0},v_{0}$, initial mesh $\mathcal{M}_{0}$, $\operatorname{tol}_{H}$, $\operatorname{tol}_{C}$, final time $T$, number of time-steps $N$, Dörfler marking threshold $\theta>0$
2:adapted meshes $(\mathcal{M}_{n})_{n=0,\ldots,N}$ and discrete solution $(U^{n})_{n=0,\ldots,N}$ such that $\|U-u\|<\operatorname{tol}$
3:procedure adapt($T,\operatorname{tol}_{H},\operatorname{tol}_{C},\Delta t,\theta,u^{0},v^{0}$)
@@
8:         set $\varepsilon_{0}^{n+1},\varepsilon_{1}^{n+1}\leftarrow\nicefrac{{2\operatorname{tol}_{H}}}{{N}}$
9:         while $\min{(\varepsilon_{0}^{n+1},\varepsilon_{1}^{n+1})}>\nicefrac{{\operatorname{tol}_{H}}}{{N}}$ do
@@
11:              compute $\varepsilon_{0}^{n+1},\varepsilon_{1}^{n+1}$triangleright store $(\eta_{i}^{K})_{K},i=0,1$
@@
16$\triangleright$ where $\mathcal{M}_{n+1}$ is in decreasing order of $\eta_{i}^{K}$
17:                       if $s\leq\theta\,(\varepsilon_{i}^{n+1})^{2}$ then
@@
28:              if $\max{(\beta_{K}^{0},\beta_{K}^{1})}<\operatorname{tol}_{c}$ then
@@
    In line 15, we use a Dörfler marking strategy for refinement [Verfurth:13:book:A-posteriori].
@@
Before starting the actual time integration, the initial uniform mesh $\mathcal{M}_{0}$ is adaptively refined to better resolve the initial conditions $u_{0}$ and $v_{0}$, resulting in the mesh $\mathcal{M}_{1}$. First, we estimate the local error indicator $\eta_{0}$ for $u_{0}$ by computing the element-wise product of the local mesh size and the $\operatorname{L}^{2}$-norm of its second spatial derivative. Elements for which $\eta_{0}(K)>\theta\max_{K}\{\eta_{0}(K)\}$ are marked for refinement and refined using a Dörfler marking strategy [Verfurth:13:book:A-posteriori]. The same procedure is applied to $v_{0}$ using a second error indicator $\eta_{1}$, which eventually yields the initial mesh $\mathcal{M}_{1}$. The initial data $u_{0}$ and $v_{0}$ are then discretized on $\mathcal{M}_{1}$ via interpolation. Finally, the first time step $U^{1}$ is computed using a Taylor expansion.
@@
1:$u^{0},v^{0},\mathcal{M}_{0}$,$\theta$
2:$U^{0},U^{1},\mathcal{M}_{1}$
3:procedure initialize($u^{0},v^{0},\mathcal{M}_{0}$,$\theta$)
@@ 4 Numerical results
Here we present a series of numerical experiments which confirm the optimal convergence rates of the LF-LTS-FEM method (2.17) with mesh change and illustrate the usefulness of the space-time adaptive algorithm from Section 3.4. First, we consider the solution of the inhomogeneous wave equation (2.1) on a predefined time-evolving mesh under successive mesh refinement to verify convergence. Then, we apply the space-time adaptive LF-FEM algorithm to compute two distinct solutions: a right-moving Gaussian pulse and a Gaussian pulse which splits into two symmetric waves traveling in opposite directions.
In all numerical experiments, we consider the one-dimensional wave equation (2.1) in $\Omega=(-10,10)$ with homogeneous Dirichlet boundary conditions, i.e. $\Gamma=\Gamma_{D}$, and $c\equiv 1$. We always use standard piecewise linear $\operatorname{H}^{1}$-conforming finite elements on a nonuniform mesh with mass-lumping in space and the leapfrog-based local time-stepping (LF-LTS) method with global time-step $\Delta t$ (without added stabilization) from Section 3.3. For the space-time adaptive LF-FEM algorithm we set the parameters $tol_{H}=20$, $tol_{C}=10^{-4}$ and $\theta=0.8$.
@@ 4.2 Traveling wave
The entire space-time evolving mesh with $h^{c}=0.11$ is shown in Fig. 5A. We observe how the fine part of the mesh $\mathcal{M}_{n}^{f}$ automatically follows the peak of the Gaussian pulse as it propagates rightward across $\Omega$. The mesh (and hence the associated FE space $\mathbbmss{V}_{n}$) changes whenever the maximum of the elliptic error indicators $\varepsilon^{n}_{0}$ and $\varepsilon^{n}_{1}$ is larger than the given tolerance $tol_{H}$ divided by the total number of time-steps. Hence the fine mesh $\mathcal{M}^{f}_{n}$ moves to the right, with increasing $n$, at the same unit wave speed as the pulse, while two subsequent meshes $\mathbbmss{V}_{n}$ and $\mathbbmss{V}_{n+1}$ always remain compatible during any mesh change, see Section 3.2. The remaining vertical lines inside the mesh are due to small dispersive waves in the discrete solution, which lead to localized refinement in regions of higher curvature.
@@ 4.3 Splitting wave
Finally, we consider a Gaussian pulse which splits into two symmetric waves traveling in opposite directions with constant speed $c\equiv 1$. The initial conditions $u_{0},v_{0}$ are set such that exact solution, centered about $x=1$ at $t=0$, is
@@
Again we apply the space-time adaptive algorithm from Section 3.4 as described in Section 4.2. We observe how the fine part of the mesh $\mathcal{M}_{n}^{f}$ automatically adapts and follows the two separating peaks of the Gaussian pulse. In particular, the initially refined single interval automatically splits into two separate refined regions, each associated with its own local time-step. That topological change in the refined part of the mesh requires no particular attention or reordering of the unknowns, as the time integration is fully explicit. The mesh (and hence the associated FE space $\mathbbmss{V}_{n}$) again changes whenever the maximum of the elliptic error indicators $\varepsilon^{n}_{0}$ and $\varepsilon^{n}_{1}$ is larger than the given tolerance $tol_{H}$ divided by the total number of time-steps. During any mesh change, two subsequent meshes $\mathbbmss{V}_{n}$ and $\mathbbmss{V}_{n+1}$ always remain compatible, as shown in Fig. 7A for $h^{c}=0.11$.
@@ 5 Concluding remarks
We have proposed a space-time adaptive strategy for the solution of the time-dependent wave equation which combines three essential ingredients for robustness and efficiency. First, the rigorous a posteriori estimates [GroteLakkisSantos:24:PosterioriErrorEstimatesWave] from Section 3.1 provide guaranteed error bounds that include all sources of error. As the error indicators are fully computable and local, they permit to decide ”on the fly” whether to accept, or recompute on an adapted mesh, the numerical solution at the current time-step. Second, by restricting mesh change to compatible meshes, as described in Section 3.2, we ensure that local mesh refinement never leads to a loss in accuracy while allowing coarsening only where appropriate to minimize the resulting inherent information loss from interpolation or projection. Third, we apply leapfrog based explicit local time-stepping [DiazGrote:09:article:Energy-Conserving, GroteMitkova:10:article:Explicit-Local, GroteMehlinSauter:18:article:Convergence, GroteMichelSauter:21:article:Stabilized] inside any locally refined region of the mesh, each associated with its own local time-step. Thus we overcome the bottleneck caused by local mesh refinement due to any overly stringent CFL stability condition while retaining the simplicity, efficiency and inherent parallelism of explicit time integration.
Our numerical experiments confirm the expected order of convergence and illustrate the usefulness of our adaptive strategy. Although it is only studied in one spatial dimension, our space-time adaptive strategy, including the a posteriori error bounds and local time-stepping approach, is dimension independent and immediately applies to two or three space dimensions. Yet even in a single space dimension, the computational effort and memory requirement are up to two orders of magnitude smaller than a standard approach on a uniform mesh. Clearly, that reduction in computational cost and memory requirement will be even more pronounced in higher dimensions, although operations related to mesh change then also become more involved. The overall gain in efficiency compared to a static, uniform space-time mesh will generally depend on the problem at hand and in particular on the size of the locally refined region relative to the entire space-time mesh. Nonetheless the reduction in the number of degrees of freedom will generally enable simulations of improved accuracy, which otherwise would be prohibitive merely due to shear problem size.
[…]

### v3

@@ 3.3 Local time-stepping
Here we briefly recall the leapfrog (LF) based local time-stepping (LF-LTS) method used for time integration during the adaptive algorithm to circumvent the bottleneck otherwise caused by local refinement on standard explicit time-stepping methods. The original LF-LTS method for the numerical solution of the second-order wave equations was proposed for homogeneous right-hand sides in [undefm] and for inhomogeneous right-hand sides in [undefz, Sect. 4.1]. Inside the locally refined region, the latter applies standard LF time-marching with a smaller time-step, which also implies evaluating the inhomogeneous right-hand side at all intermediate times. Although we focus on second-order LTS-LF here, we note that higher-order versions are available [undefm]. For $f=0$ they also conserve (a discrete version of) the energy. Optimal convergence rates for the LF-LTS method from [undefm] with $p$ local time steps were derived for a conforming FEM discretization, albeit under a sub-optimal CFL condition where $\Delta t$ in fact depends on the smallest elements in the mesh [undefw].
[…]
Stabilization was also recently introduced into the original LTS-LF(0)(0) algorithm from [undefz] for nonzero source terms, which again led to optimal $L^{2}$-convergence rates under a CFL condition independent of the coarse-to-fine mesh ratio [undefx]. Similar optimal $L^{2}$-convergence rates for a somewhat simpler ”split-LFC” variant, which omits intermediate source evaluations inside the refined region, were recently proved in [undefh].
@@
[…]
@@ 3.4 Space-time adaptive algorithm
Given a fixed ”coarse” FE mesh, $\mathcal{M}_{0}$, which discretizes the computational domain $\Omega$, we first choose a constant global time-step $\Delta t$ which satisfies the CFL stability condition imposed by $\mathcal{M}_{0}$ on the (standard) leapfrog method. Next, the mesh is adapted to the initial conditions $u_{0},v_{0}$ - see algorithm 2 – which yields the initial mesh $\mathcal{M}_{1}$ used for time integration. Now, during each time step, we proceed as follows. First, we tentatively set $\mathcal{M}_{n+1}$ to the current mesh $\mathcal{M}_{n}$ and compute $U^{n+1}$ using the LTS-LF method inside locally refined regions and the standard LF method elsewhere. Next, we evaluate the elliptic error indicators $\varepsilon_{0}$ and $\varepsilon_{1}$ in (3.6) from $U^{n+1}$ and refine $\mathcal{M}_{n+1}$ accordingly, if needed. We then transfer $U^{n-1}$ and $U^{n}$ to the new finite element space $\mathbbmss{V}_{n+1}$, associated with $\mathcal{M}_{n+1}$ via interpolation or $\operatorname{L}^{2}$-projection, and repeat the computation of $U^{n+1}$. This process – refining $\mathcal{M}_{n+1}$ and recomputing $U^{n+1}$ – is iterated until $\varepsilon_{0}$ or $\varepsilon_{1}$ falls below the threshold $\operatorname{tol}_{\text{H}}/{N}$. At that point, $U^{n+1}$ is accepted on the current mesh $\mathcal{M}_{n+1}$. Now, the algorithm enters a coarsening step: for each $K\in\mathcal{M}_{n+1}$, we evaluate the pre-indicators $\beta_{K}$ from (3.18) and mark those elements with $\beta_{K}\leq\operatorname{tol}_{\text{C}}$. After coarsening, both $U^{n}$ and $U^{n+1}$ are transfered to the new, coarsened space $\mathbbmss{V}_{n+1}$. The entire procedure is repeated until the final time $T$ is reached.
@@
1:initial conditions $u_{0},v_{0}$, initial mesh $\mathcal{M}_{0}$, $\operatorname{tol}_{\text{H}}$, $\operatorname{tol}_{\text{C}}$, final time $T$, number of time-steps NN, Dörfler marking threshold $\theta>0$
2:adapted meshes $(\mathcal{M}_{n})_{n=0,\ldots,N}$ and discrete solution $(U^{n})_{n=0,\ldots,N}$ such that $\|U-u\|<\operatorname{tol}_{\text{}}$
3:procedure adapt($T,\operatorname{tol}_{\text{H}},\operatorname{tol}_{\text{C}},\Delta t,\theta,u^{0},v^{0}$)
@@
8:   set $\varepsilon_{i}^{n+1},\mu_{j}^{n+1},\alpha_{i}^{n+1},\vartheta_{i}^{n+1},\delta^{n+1}\leftarrow\nicefrac{{2\operatorname{tol}_{\text{H}}}}{{N}}$ $\triangleright$ $i=0,1,j=0,1,2$
9:   while $\begin{aligned} \min(\varepsilon_{i}^{n+1},\mu_{j}^{n+1},\alpha_{i}^{n+1},\vartheta_{i}^{n+1},\delta^{n+1})>\nicefrac{{\operatorname{tol}_{\text{H}}}}{{N}}\end{aligned}$ do
@@
11:     compute $\varepsilon_{i}^{n+1},\mu_{j}^{n+1},\alpha_{i}^{n+1},\vartheta_{i}^{n+1},\delta^{n+1}$ $\triangleright$ store $(\eta_{i}^{K})_{K},i=0,1$
@@
[…]
17:        if $s\leq\theta\,((\varepsilon_{i}^{n+1})^{2}+(\mu_{j}^{n+1})^{2}+(\alpha_{i}^{n+1})^{2}+(\vartheta_{i}^{n+1})^{2}+(\delta^{n+1})^{2})$ then
@@
28:     if $\max{(\beta_{K}^{0},\beta_{K}^{1})}<\operatorname{tol}_{\text{c}}$ then
@@
    In line 15, we use a Dörfler marking strategy for refinement [undefan].
@@
  7. 7.
    In line 8, the leapfrog based explicit local time-stepping method from Section 3.3 is used with a time-step ratio $p$ given by the local mesh size ratio. Thus for any refined region we use the value of $p$ determined by the ratio of the coarse to the smallest mesh size in that subregion; note that $p$ may vary from one refined subregion to another. In situations where the refined region itself contained yet another sub-region of much smaller mesh-size, a hierarchical multi-level approach could be used instead for even higher efficiency [undefn, undefai].
Before starting the actual time integration, the initial uniform mesh $\mathcal{M}_{0}$ is adaptively refined to better resolve the initial conditions $u_{0}$ and $v_{0}$, resulting in the mesh $\mathcal{M}_{1}$. First, we estimate the local error indicator $\eta_{0}$ for $u_{0}$ by computing the element-wise product of the local mesh size and the $\operatorname{L}^{2}$-norm of its second spatial derivative. Elements for which $\eta_{0}(K)>\theta\max_{K}\{\eta_{0}(K)\}$ are marked for refinement and refined using a Dörfler marking strategy [undefan]. The same procedure is applied to $v_{0}$ using a second error indicator $\eta_{1}$, which eventually yields the initial mesh $\mathcal{M}_{1}$. The initial data $u_{0}$ and $v_{0}$ are then discretized on $\mathcal{M}_{1}$ via interpolation. Finally, the first time step $U^{1}$ is computed using a Taylor expansion.
@@
[…]
3:procedure initialize($u^{0},v^{0},\mathcal{M}_{0}$$\theta$)
@@ 4 Numerical results
Here we present a series of numerical experiments that confirm the optimal convergence rates of the LF–LTS–FEM method (2.17) with mesh change and demonstrate the effectiveness of the space–time adaptive algorithm from Section 3.4. First, we consider the solution of the inhomogeneous wave equation (2.1) on a predefined, time-evolving mesh under successive mesh refinement to verify convergence. Next, we apply the space–time adaptive LF–FEM algorithm to compute two one-dimensional test cases: a right-moving Gaussian pulse and a Gaussian pulse that splits into two symmetric waves traveling in opposite directions. Finally, we consider a two-dimensional example, where a Gaussian pulse initially centered at $(0.4,0.6)$ propagates across an L-shaped domain.
In all one-dimensional experiments, we solve the wave equation (2.1) in $\Omega=(-10,10)$ with homogeneous Dirichlet boundary conditions, i.e. $\Gamma=\Gamma_{D}$, and wave speed $c\equiv 1$. We use standard piecewise linear $H^{1}$-conforming finite elements on nonuniform meshes with mass-lumping in space and the leapfrog-based local time-stepping (LF-LTS) method with global time step $\Delta t$ (with added stabilization $\nu=0.01$) from Section 3.3.
Since the entire time-stepping procedure is fully explicit—no linear systems are ever solved—the computational complexity scales linearly with the total number of degrees of freedom in both space and time. For the one-dimensional adaptive LF–FEM experiments we choose the parameters $\operatorname{tol}_{\text{H}}=20$, $\operatorname{tol}_{\text{C}}=10^{-4}$, and $\theta=0.8$ whereas in the two-dimensional experiment in Section 4.2 we set $\operatorname{tol}_{\text{C}}=0.01$.
@@ 4.2 Traveling wave
The entire space-time evolving mesh with $h^{c}=0.11$ is shown in Fig. 5A. We observe how the fine part of the mesh $\mathcal{M}_{n}^{f}$ automatically follows the peak of the Gaussian pulse as it propagates rightward across $\Omega$. The mesh (and hence the associated FE space $\mathbbmss{V}_{n}$) changes whenever the maximum of the elliptic error indicators $\varepsilon^{n}_{0}$ and $\varepsilon^{n}_{1}$ is larger than the given tolerance $\operatorname{tol}_{\text{H}}$ divided by the total number of time-steps. Hence the fine mesh $\mathcal{M}^{f}_{n}$ moves to the right, with increasing $n$, at the same unit wave speed as the pulse, while two subsequent meshes $\mathbbmss{V}_{n}$ and $\mathbbmss{V}_{n+1}$ always remain compatible during any mesh change, see Section 3.2. The remaining vertical lines inside the mesh are due to small dispersive waves in the discrete solution, which lead to localized refinement in regions of higher curvature.
@@ 4.3 Splitting wave
Next, we consider a Gaussian pulse which splits into two symmetric waves traveling in opposite directions with constant speed $c\equiv 1$. The initial conditions $u_{0},v_{0}$ are set such that exact solution, centered about $x=1$ at $t=0$, is
@@
Again we apply the space-time adaptive algorithm from Section 3.4 as described in Section 4.2. We observe how the fine part of the mesh $\mathcal{M}_{n}^{f}$ automatically adapts and follows the two separating peaks of the Gaussian pulse. In particular, the initially refined single interval automatically splits into two separate refined regions, each associated with its own local time-step. That topological change in the refined part of the mesh requires no particular attention or reordering of the unknowns, as the time integration is fully explicit. The mesh (and hence the associated FE space $\mathbbmss{V}_{n}$) again changes whenever the maximum of the elliptic error indicators $\varepsilon^{n}_{0}$ and $\varepsilon^{n}_{1}$ is larger than the given tolerance $\operatorname{tol}_{\text{H}}$ divided by the total number of time-steps. During any mesh change, two subsequent meshes $\mathbbmss{V}_{n}$ and $\mathbbmss{V}_{n+1}$ always remain compatible, as shown in Fig. 7A for $h^{c}=0.11$.
@@
4.4 L-shaped domain
Finally, we consider a Gaussian pulse which propagates with constant wave speed $c\equiv 1$ across an L-shaped domain. The initial data is chosen such that at $t=0$ the solution is a Gaussian centered at (x,y)=(0.4, 0.6)(x,y)=(0.4,\,0.6):
$u_{0}(x,y)=e^{-600\left((x-0.4)^{2}+(y-0.6)^{2}\right)},$ (4.3)
while the initial velocity $v_{0}\equiv 0$. At all boundaries we impose homogeneous Dirichlet boundary conditions and choose a uniform initial mesh $\mathcal{M}_{0}$ with $h=0.08$. Again we apply the space–time adaptive algorithm from Section 3.4, following the set-up described in Section 4.2. The refinement and coarsening are performed using the iFEM package [undefk], combined with our stabilized LF–LTS method for time integration with a fixed stabilization parameter $\nu=0.01$.
As shown in Figures 9A and 9C, the refined region of the mesh $\mathcal{M}_{n}^{f}$ automatically adapts and tracks the propagating wave front. The corresponding adaptive meshes are shown in Figures 9B and 9D. In Fig. 10A, we display the elliptic error indicator $\varepsilon_{0}^{n}$ (3.6) from Section 3.1 without accumulation over time. Finally, in Fig. 10B, we show the relative memory requirement of the adaptive LF-FEM by computing the ratio of the number of dof’s (FE nodes in the mesh) for the adaptive LF-FEM algorithm vs. that using a uniform mesh.
9A L-shaped domain: Numerical solution at time $t=0.1$.
9B Space-time adaptive mesh at time $t=0.1$.
9C Numerical solution at time $t=0.3$.
9D Space-time adaptive mesh at time $t=0.3$.
10A L-shaped domain: Elliptic error indicator $\varepsilon^{n}_{0}$ in (3.6) vs. time without accumulation.
10B Relative memory w.r.t. uniform refinement.
@@ 5 Concluding remarks
We have proposed a space-time adaptive strategy for the solution of the time-dependent wave equation which combines three essential ingredients for robustness and efficiency. First, the rigorous a posteriori estimates [undefv] from Section 3.1 provide guaranteed error bounds that include all sources of error. As the error indicators are fully computable and local, they permit to decide ”on the fly” whether to accept, or recompute on an adapted mesh, the numerical solution at the current time-step. Second, by restricting mesh change to compatible meshes, as described in Section 3.2, we ensure that local mesh refinement never leads to a loss in accuracy while allowing coarsening only where appropriate to minimize the resulting inherent information loss from interpolation or projection. Third, we apply leapfrog based explicit local time-stepping [undefm, undefz, undefw, undefy] inside any locally refined region of the mesh, each associated with its own local time-step. Thus we overcome the bottleneck caused by local mesh refinement due to any overly stringent CFL stability condition while retaining the simplicity, efficiency and inherent parallelism of explicit time integration.
Although our adaptive algorithm relies on a fixed underlying coarse mesh and a constant global time-step, it nonetheless automatically adapts both the mesh and the time-step to capture locally the waves’ smaller scale features as they propagate across the computational domain. While the increased flexibility of a varying global time-step might be all too tempting, one must keep in mind that the standard leapfrog method with varying time-step is prone to instability [undefak].
Our numerical experiments confirm the expected order of convergence and illustrate the usefulness of our adaptive strategy. Although it is only studied in one and two space dimensions, our space-time adaptive strategy, including the a posteriori error bounds and local time-stepping approach, are dimension independent and immediately apply three space dimensions, too. Yet even in a single space dimension, the computational effort and memory requirement are up to two orders of magnitude smaller than a standard approach on a uniform mesh. Clearly, that reduction in computational cost and memory requirement will be even more pronounced in higher dimensions, although operations related to mesh change then also become more involved. The overall gain in efficiency compared to a static, uniform space-time mesh will generally depend on the problem at hand and in particular on the size of the locally refined region relative to the entire space-time mesh. Nonetheless the reduction in the number of degrees of freedom will generally enable simulations of improved accuracy, which otherwise would be prohibitive merely due to sheer problem size.
Acknowledgements
This work was supported by the Swiss National Science Foundation under grant SNF 200020-188583.